# Learning Patterns from Biological Networks: A Compounded Burr Probability Model

### v2

@@ Abstract
Complex biological networks, encompassing metabolic reactions, gene interactions, and protein-protein interactions, often exhibit scale-free characteristics with power-law degree distributions. However, empirical evidence reveals significant deviations from ideal power-law fits, necessitating more flexible and accurate modeling approaches. To address this challenge, we introduce a novel Compounded Burr (CBurr) distribution, a novel probability model derived from the Burr family, designed to capture the intricate structural properties of biological networks. We rigorously establish its statistical properties, including moment analysis, hazard functions, and tail behavior, and provide a robust parameter estimation framework using the maximum likelihood method. The CBurr distribution is broadly applicable to networks with fat-tailed degree distributions, making it highly relevant for modeling biological, social, and technological networks. To validate its efficacy, we conduct an extensive empirical study on large-scale biological network datasets, demonstrating that CBurr consistently outperforms conventional power-law and alternative heavy-tailed models in fitting the entire range of node degree distributions. Our proposed CBurr probability distribution holds great promise for accurately capturing the complex nature of biological networks and advancing our understanding of their underlying mechanisms.
@@ I Introduction
Biological networks that follow a scale-free structure exhibit a distinctive pattern of connectivity, where a few nodes are highly interconnected (known as hubs), while the majority of nodes have relatively lower degrees of connectivity Barabási and Albert (1999); Broido and Clauset (2019); Chattopadhyay et al. (2021a). These networks provide valuable insights into the resilience and efficiency of biological systems, serving as a foundation for studying intricate biological processes Wuchty (2001); Arita (2005); Khanin and Wit (2006); Wuchty, Ravasz, and Barabási (2006). Lately, there has been a surge in curiosity regarding the exploration of biological networks in real-world settings, including metabolic networks Jeong et al. (2000); Fell and Wagner (2000); Ma and Zeng (2003), protein-protein interaction networks Jeong et al. (2000); Uetz et al. (2000); Schwikowski, Uetz, and Fields (2000); Maslov and Sneppen (2002); Rual et al. (2005), protein domain networks Rzhetsky and Gomez (2001); Wuchty (2001), gene interactions Tong et al. (2004), and genetic regulatory networks Lee et al. (2002); Farkas et al. (2003); Luscombe et al. (2004). These networks are considered real-world complex systems and have been predominantly modeled using graph theory Haggarty, Clemons, and Schreiber (2003); Dorogovtsev and Mendes (2003). Metabolic networks have been extensively studied, providing insights into the interconnected biochemical reactions that drive cellular processes Ma and Zeng (2003); Jeong et al. (2000); Fell and Wagner (2000). Protein-protein interaction networks have been explored, shedding light on the physical interactions between proteins and their functional implications Jeong et al. (2000); Maslov and Sneppen (2002); Uetz et al. (2000); Rual et al. (2005); Schwikowski, Uetz, and Fields (2000). Protein domain networks have been investigated, revealing the modular organization of proteins and their functional domains Rzhetsky and Gomez (2001); Wuchty (2001). Gene interactions and genetic regulatory networks have been analyzed, elucidating the complex regulatory mechanisms that govern gene expression and cellular processes Tong et al. (2004); Lee et al. (2002); Farkas et al. (2003); Luscombe et al. (2004).
Biological networks provide a valuable framework for modeling the structure of biological systems Lima-Mendez and Van Helden (2009); Miranda et al. (2016); Saucan, Samal, and Jost (2021). In these networks, nodes represent biological elements, and edges connect elements that interact with each other Brede (2012); Barabási (2013); Albert and Barabási (2002). The number of edges connected to each node, known as its node degree, is categorized as in-degree for incoming edges and out-degree for outgoing edges. Studying the distribution of node degrees is essential for understanding the connectivity patterns within a network and identifying hubs—nodes with a high number of connections Artime et al. (2024). Degree distribution indicates whether the network has a homogeneous or heterogeneous structure in terms of node connectivity. Understanding these structural patterns is essential for modeling how networks evolve over time, grow or shrink, and how new nodes and edges are likely to form. By utilizing graph theory and analyzing the degree distribution, researchers gain a deeper understanding of the functional and structural characteristics of complex biological systems Barabási and Albert (1999); Newman (2003); Clote (2020). Modeling biological systems through biological networks offers a powerful approach to studying the intricate interactions and behaviors within these systems, ultimately contributing to a deeper understanding of the underlying mechanisms Barabási and Albert (1999); Newman (2003). Cutting-edge advancements in deep learning techniques have provided a formidable approach to deciphering intricate patterns and extracting valuable insights from intricate biological networks Muzio, O’Bray, and Borgwardt (2021); Jin et al. (2021); Pan et al. (2022); Yang et al. (2022); Ma et al. (2023).
Standard existing distributions like power law, exponential, log-normal, Poisson, etc., have been widely used to model degree distributions in different types of networks. However, these distributions may not always accurately capture the intricacies and nuances of real-world complex networks, which often exhibit heterogeneity, assortativity, and other complex patterns. Khanin and Wit Khanin and Wit (2006) highlighted that gene interaction networks determined by synthetic lethal interactions Tong et al. (2004), metabolic interaction networks Tong et al. (2004), and protein-protein interaction networks Ito et al. (2000); Schwikowski, Uetz, and Fields (2000) do not follow a power-law distribution, although some of these networks may exhibit a truncated power-law distribution. Clauset et al. Clauset, Shalizi, and Newman (2009) analyzed various real-world datasets and demonstrated that while some datasets supported the power-law claim, others did not, using maximum-likelihood estimation and statistical tests for goodness-of-fit, utilizing Kolmogorov-Smirnov statistic and likelihood ratios. Salem Salem (2018) discussed the impact of disturbances in biological networks on disease susceptibility in living cells. A study by Goldstein and Morris Goldstein, Morris, and Yen (2004) highlighted the importance of evaluating the goodness-of-fit of the distribution and discovered that the power-law distribution is not an adequate fit for certain cases. On the other hand, Clauset et al. Clauset, Shalizi, and Newman (2009) demonstrated various parameter estimation methods and concluded that while the power-law distribution fits some datasets well, it fails to represent other datasets accurately. Broido and Clauset Broido and Clauset (2019) argued that strictly scale-free networks are rare and introduced notions of weak or strong scale-free networks. Voitalov et al. Voitalov et al. (2019) demonstrated a close relationship between the scale-free property and the Pareto type-II or Lomax distribution. Some recent studies have demonstrated the use of Lomax Chattopadhyay et al. (2021b), Burr Chakraborty et al. (2022), and other heavy-tailed distributions Chakraborty, Das, and Chattopadhyay (2022) for modeling real-world complex networks. Broido and Clauset Broido and Clauset (2019) further noted that recent complex network datasets deviate from the power-law distribution, supporting earlier claims by Stumpf, Wiuf, and May (2005). In 2020, Clote Clote (2020) presented an efficient algorithm for computing the connectivity density function of homopolymer secondary structures, addressing the computational challenges posed by the exponential size of such networks. The obtained hypothesis tests strongly indicated that homopolymer RNA secondary structure networks are not scale-free. Introducing a new degree distribution allows us to provide a better fit to empirical data, thereby improving the accuracy of network analysis models. Additionally, a novel distribution may uncover important hidden patterns or phenomena within networks. It can also critically evaluate existing assumptions about network structure and behavior. Therefore, by exploring different parameterizations and characteristics of the new distribution, we can reveal hidden aspects of network organization and behavior that were not apparent through the existing models. This can lead to new theoretical discoveries and practical applications across various disciplines, from sociology and biology to computer science and economics. Hence, introducing a new degree distribution enhances the robustness and reliability of network analysis methods, facilitating more accurate predictions and insights into real-world systems.
Many real-world complex networks, including biological networks, do not strictly adhere to power-law distributions; instead, they may exhibit characteristics such as truncated tails, multimodality, or deviations from pure scale-free behavior. In such cases, the Burr distribution can provide a better empirical fit, allowing us to model network data more accurately and extract more reliable insights into network structure and behavior Burr (1942). The Burr distribution is highly flexible and versatile, capable of modeling a wide range of shapes and heavy-tail behaviors. Due to its flexibility, it might be able to fit network degree distributions with different levels of heterogeneity and heavy tails that are common in biological networks. The Burr distribution has well-defined statistical properties, including moments, cumulative distribution function, and tail behavior indices (e.g., shape parameters), which are essential for quantitative analysis in network science. These properties may allow us to derive meaningful interpretations about network characteristics such as centralization, assortativity, and resilience based on the distribution of node degrees. This paper introduces a novel family of distributions with the Burr degree distribution as the baseline. Recently, Burr distribution with nonlinear exponent Chakraborty, Das, and Chattopadhyay (2022), Marhsall-Olkin generalization of Burr (Burr MO Jayakumar and Mathew (2008)), and exponentiated Burr distribution Ahmed et al. (2021) have shown excellent performance to better fit the large-scale heavy-tailed networks better than the popularly used power law and other probability models. However, for a few cases, it cannot cover the entire distribution pattern of complex large-scale networks Chakraborty et al. (2022). Thus, a modification or a more accurate modelling of the degree distribution is required. The proposed Compounded Burr (CBurr) distribution is a member of a specific family of continuous probability distributions, which introduces a new way of incorporating a parameter to expand the Burr XII distribution. The proposed distribution aims to improve the fitting of biological network data across the entire range, outperforming commonly used power-law distributions such as Pareto, Lomax, Burr distributions, and some other distributions in most biological networks.
Our proposal increases the flexibility of the model, allowing it to better capture the nuances of the degree distributions in complex networks, particularly when the network exhibits a mixture of different subpopulations with varying connectivity patterns. Real-world complex networks, such as social networks, biological networks, and technological networks, often show multi-scale behavior in their degree distributions and, hence, do not conform perfectly to a single distribution form Wang, Ma, and Yao (2021). Thus, the compounded Burr distribution, by blending two distributions, can provide a better fit and is well-suited to capture these behaviors, as it can model both the heavy-tailed nature of the distribution and the varying scales of connectivity present within different parts of the network. Therefore, it may lead to a more accurate modeling of the network’s structure, which is crucial for understanding its dynamics and resilience. We have demonstrated the general properties of this new distribution family, including maximum likelihood estimation. Furthermore, the proposed CBurr distribution model is applied to comprehensively analyze large-scale biological network data. This analysis aims to demonstrate the effectiveness of the proposed distribution in capturing the intricate connectivity and behavior exhibited by biological networks, surpassing the limitations of power-law models. In essence, the proposed model enables effective modeling of the degree distribution of complex biological networks across the entire range of data, eliminating the need to discard lower-degree nodes. The practical application of this novel distribution family is demonstrated using large-scale heavy-tailed biological network datasets surpassing the performance of commonly utilized heavy-tailed distributions. Overall, the proposed CBurr distribution for modeling the degree distribution of biological networks aims to represent network connectivity accurately, improve biological interpretation, network analysis, and prediction, and develop new statistical models.
@@
The structure of the remaining paper is as follows: Section II discusses a motivating example from biological networks, and Section III presents the genesis of the proposed family of distribution and its statistical properties to provide a better understanding of the proposed model. Section IV introduces the Compounded Burr (CBurr) distribution, including its statistical properties such as parameter estimation. In Section V, we present experimental results and analyze them in the context of various real-world complex networks. Finally, Section VI concludes the paper with a concise summary and discussion.
II Motivating Examples
@@
Figure 1: Plots of the node degree distribution of bio-Dmela network in the Original scale (left) and in the Log-Log scale (right). It is clear from the plot (right) that power law distribution does not provide a ‘good’ fitting for the bio-Dmela data, and there is a need for heavy-tailed probability distribution to provide a ‘good’ nonlinear fitting.
@@
One of the key structural features in analyzing real-world complex networks is the degree distribution. Degree distribution of nodes, especially a power-law degree distribution, has been regarded as one of the most significant structural characteristics of social, information, and biological networks. Empirical studies on real-world network patterns suggest that their degree distributions follow a single power-law, often called a scale-free network. The main reason behind the emergence of this phenomenon is the mechanism of preferential attachment, which states that in a growing network, a node with a higher degree is more likely to receive new links. In 1999, Barabási Barabási and Albert (1999) modeled the degree distribution of the World Wide Web (WWW) using a single power-law. Since then, this structural pattern has been extensively studied across various real-world networks, including metabolic, gene regulatory, collaboration, communication, and social networks. These networks are inherently dynamic, as their vertices and/or edges continue to grow over time. This dynamic nature can be effectively captured by studying the underlying pattern of node degree distribution.
@@
However, upon closer examination of the entire network data, this single power-law distribution often fails to represent the data accurately because of the presence of large fluctuations and sparsity in the upper and lower tails of the degree distribution. If the degree distribution followed a single power law, then a single straight line would be enough to fit all the data points completely when viewed on the log-log scale. However, Fig. 1 (right) clearly demonstrates that this is not the case. We analyzed a bio-Dmela heavy-tailed network Singh, Xu, and Berger (2008a) and visualized the node degree distribution on both the original scale and the log-log scale (see Fig. 1). When attempting to fit a single power-law model to the entire bio-Dmela network data, we observed that the blue straight line (see Fig. 1 (right)) representing the power-law fit failed to align with many data points, particularly those corresponding to lower-degree nodes, indicating a poor fit in those regions. This suggests that when modeling with a single power law, greater emphasis is placed on higher-degree nodes, while many lower-degree nodes, potentially containing valuable information about the network, are overlooked or disregarded. Finally, our observation from Fig. 1 (right) reveals that the node degree distribution of the bio-Dmela network, when plotted on a log-log scale, exhibits nonlinearity (as shown by the yellow curve in Fig. 1 (right)) upon closer inspection. This behavior differs from the pattern observed when a single power-law model is applied and has also been noted in many other biological networks. Therefore, when fitting the degree distribution, it is crucial to consider all nodes, including both lower-degree and higher-degree nodes, to accurately capture the true pattern of the network’s degree distribution. This is why more research is being done to find a model that can accurately describe the main features of heavy-tailed and long-tailed behavior in the entire degree distribution of real-world complex networks. This article contributes to that effort by introducing a novel heavy-tailed compounded Burr distribution, a variant of the widely used heavy-tailed Burr distribution for effectively capturing the structural characteristic of the node degree distribution of real world networks.

### v3

# A Compounded Burr Probability Distribution for Fitting Heavy-Tailed Data with Applications to Biological Networks

@@ Abstract
Complex biological networks, encompassing metabolic pathways, gene regulatory systems, and protein–protein interaction networks, often exhibit scale-free structures characterized by heavy-tailed degree distributions. However, empirical studies reveal significant deviations from ideal power-law behavior, underscoring the need for more flexible and accurate probabilistic models. In this work, we propose the Compounded Burr (CBurr) distribution, a novel four-parameter family derived by compounding the Burr distribution with a discrete mixing process. This model is specifically designed to capture both the body and tail behavior of real-world network degree distributions with applications to biological networks. We rigorously derive its statistical properties, including moments, hazard and risk functions, and tail behavior, and develop an efficient maximum likelihood estimation framework. The CBurr model demonstrates broad applicability to networks with complex connectivity patterns, particularly in biological, social, and technological domains. Extensive experiments on large-scale biological network datasets show that CBurr consistently outperforms classical power-law, log-normal, and other heavy-tailed models across the full degree spectrum. By providing a statistically grounded and interpretable framework, the CBurr model enhances our ability to characterize the structural heterogeneity of biological networks.
@@ I Introduction
Biological networks often exhibit a scale-free topology, where a small number of nodes act as highly connected hubs, while the majority possess relatively few connections Barabási and Albert (1999); Broido and Clauset (2019); Chattopadhyay et al. (2021a). This heterogeneous structure provides a crucial lens into the organizational principles, efficiency, and resilience of biological systems Wuchty (2001); Arita (2005); Khanin and Wit (2006); Wuchty, Ravasz, and Barabási (2006). Driven by this, considerable efforts have been made to model real-world biological networks such as metabolic networks Jeong et al. (2000); Fell and Wagner (2000); Ma and Zeng (2003), protein-protein interaction networks Jeong et al. (2000); Uetz et al. (2000); Schwikowski, Uetz, and Fields (2000); Maslov and Sneppen (2002); Rual et al. (2005), protein domain networks Rzhetsky and Gomez (2001); Wuchty (2001), gene interaction maps Tong et al. (2004), and genetic regulatory networks Lee et al. (2002); Farkas et al. (2003); Luscombe et al. (2004). These networks represent real-world complex systems and have been widely analyzed through the lens of graph theory Haggarty, Clemons, and Schreiber (2003); Dorogovtsev and Mendes (2003). Biological networks serve as a valuable framework for characterizing the intricate structure of biological systems Lima-Mendez and Van Helden (2009); Miranda et al. (2016); Saucan, Samal, and Jost (2021). In such networks, nodes represent biological entities (e.g., genes, proteins, metabolites), and edges capture interactions or functional relationships between them Brede (2012); Barabási (2013); Albert and Barabási (2002). A node’s connectivity—its degree—is categorized into in-degree (incoming edges) and out-degree (outgoing edges), and understanding the statistical distribution of degrees. This is critical for identifying central nodes and capturing global structural properties Artime et al. (2024).
The degree distribution plays a pivotal role in distinguishing between homogeneous and heterogeneous topologies, guiding how networks evolve, adapt, and reorganize. Analytical tools based on graph theory and degree distribution allow researchers to decode the structural and functional characteristics of complex biological systems Barabási and Albert (1999); Newman (2003); Clote (2020). The advent of deep learning has further augmented our ability to extract hidden patterns from these systems, enhancing biological interpretation and predictive modeling Muzio, O’Bray, and Borgwardt (2021); Jin et al. (2021); Pan et al. (2022); Yang et al. (2022); Ma et al. (2023). Although classical probability models like power-law, exponential, log-normal, and Poisson distributions have been extensively applied to model degree distributions, these approaches often fall short in accurately capturing the nuances and heterogeneity inherent in heavy-tailed biological networks. For instance, Khanin and Wit Khanin and Wit (2006) found that many biological networks, such as gene interaction networks defined by synthetic lethal interactions Tong et al. (2004), and metabolic and protein interaction networks Ito et al. (2000); Schwikowski, Uetz, and Fields (2000), deviate from pure power-law behavior. This has been echoed by Clauset et al. Clauset, Shalizi, and Newman (2009), who showed that while some datasets conform to a power-law, others do not, based on rigorous goodness-of-fit tests and maximum likelihood estimations.
Similar critiques have emerged in subsequent works Goldstein, Morris, and Yen (2004); Salem (2018), highlighting the need for distributions that better capture complex real-world network topologies. Broido and Clauset Broido and Clauset (2019) underscored the rarity of strictly scale-free networks and introduced nuanced classifications such as weak or strong scale-free types. Clote Clote (2020) provided an efficient algorithm for estimating the degree distribution in RNA secondary structure networks, revealing that these structures do not conform to power-law models. These limitations in existing distributions necessitate the development of new models capable of accurately capturing the entire degree spectrum of biological networks. Voitalov et al. Voitalov et al. (2019) also demonstrated that Pareto type-II (Lomax) distributions may better approximate scale-free behavior. Recent efforts have introduced more flexible alternatives such as the modified Lomax Chattopadhyay et al. (2021b), Burr Chakraborty et al. (2022), and other heavy-tailed models Chakraborty, Das, and Chattopadhyay (2022). Still, even these enhanced models sometimes fail to represent the full range of connectivity, particularly among low-degree nodes for heavy-tailed complex networks Stumpf, Wiuf, and May (2005). The Burr distribution Burr (1942), known for its versatility and capacity to model heavy-tailed behaviors, offers a promising foundation. It can flexibly accommodate diverse topologies and levels of heterogeneity observed in biological networks. Theoretical extensions such as the exponentiated Burr Ahmed et al. (2021) and its nonlinear variants Chakraborty, Das, and Chattopadhyay (2022), and Marshall-Olkin generalized Burr (Burr-MO) Jayakumar and Mathew (2008) have demonstrated superior performance across a variety of complex network settings. However, as shown in Chakraborty et al. (2022), even these models sometimes fall short of fully capturing the distributional complexities of large-scale heavy-tailed biological networks.
Motivated by the above discussion, this study proposes a novel Compounded Burr (CBurr) distribution to model the degree distributions of biological networks more accurately. By compounding the Burr distribution with a Poisson structure (introducing a Poison-shifted parameter), the proposed distribution enhances flexibility and better reflects empirical network characteristics — especially in cases where networks exhibit heavy truncated tails. The CBurr model introduces a new parameterization that captures both the heavy-tailed nature and heterogeneity observed in real data. Our empirical evaluation demonstrates that CBurr consistently outperforms classical distributions across various biological network datasets, without the need to discard low-degree nodes. Furthermore, we establish its theoretical properties, including estimation via maximum likelihood, making it practically useful.
Real-world networks such as those in biology, sociology, and technology often reflect multiscale behaviors that defy strict adherence to classical distributional assumptions Wang, Ma, and Yao (2021). Hence, the CBurr distribution, by blending multiple distributional behaviors, offers a robust alternative for modeling the complex and diverse architectures of real-world networks. Unlike social or technological networks, where preferential attachment may continue relatively unchecked, biological systems are governed by physical, functional, and evolutionary constraints. These constraints motivate building Poisson-Shifted Burr (we named it CBurr) model, which incorporates an initial heavy tail followed by a softened decay, aligning well with biological realism. Moreover, the study of risk and residual life functions in this context can directly inform biological interpretations, such as estimating the likelihood of further molecular interactions or identifying saturation effects in complex signaling pathways. While the model is broadly applicable, starting with biological networks ensures a tight linkage between statistical properties and biological meaning, and paves the way for future generalizations to other domains.
@@
The structure of the remaining paper is as follows: Section II discusses a motivating example from biological networks, and Section III presents the genesis of the proposed family of distributions and its statistical properties to provide a better understanding of the proposed model. Section IV introduces the Compounded Burr (CBurr) distribution, including its statistical properties such as parameter estimation. In Section V, we present experimental results and analyze them in the context of various real-world complex networks. Finally, Section VI concludes the paper with a concise summary and discussion.
II Motivating Example
@@
Figure 1: Plots of the node degree distribution of bio-Dmela network in the Original scale (left) and in the Log-Log scale (right). It is clear from the plot (right) that a power-law distribution does not provide a ‘good’ fitting for the bio-Dmela data, and there is a need for a more flexible heavy-tailed probability distribution to provide a ‘good’ nonlinear fitting.
@@
One of the key structural features in analyzing real-world complex networks is the degree distribution. Degree distribution of nodes is one of the most significant structural characteristics of social, information, and biological networks. Previous empirical studies on real-world network patterns suggest that their degree distributions follow a single power-law, often called a scale-free network Newman (2003). The main reason behind the emergence of this phenomenon is the mechanism of preferential attachment, which states that in a growing network, a node with a higher degree is more likely to receive new links. In 1999, Barabási Barabási and Albert (1999) modeled the degree distribution of the World Wide Web (WWW) using a single power-law. Since then, this structural pattern has been extensively studied across various real-world networks, including metabolic, gene regulatory, collaboration, communication, and social networks Brede (2012). These networks are inherently dynamic, as their vertices and/or edges continue to grow over time. This dynamic nature can be effectively captured by studying the underlying pattern of node degree distribution. However, upon closer examination of the entire network data, this single power-law distribution often fails to accurately represent the data due to the presence of large fluctuations and sparsity in the upper and lower tails of the degree distribution. If the degree distribution followed a single power-law, then a single straight line would be enough to fit all the data points completely when viewed on the log-log scale.
@@
For illustration, we consider bio-dmela (Drosophila melanogaster) PPI (Protein-Protein Interaction) networks, which are characterized by their heavy-tailed behaviour Zhang and Zhang (2019). These networks are typically highly skewed, with many low-degree nodes and few hubs (high-degree nodes). In particular, bio-dmela are known to follow sub-exponential or power-law–like distributions, but often deviate from a pure power-law due to biological constraints or network truncations Broido and Clauset (2019). We analyze a bio-Dmela heavy-tailed network Singh, Xu, and Berger (2008a) and visualize the node degree distribution on both the original scale and the log-log scale (see Fig. 1) as a motivating example. When attempting to fit a single power-law model to the entire bio-Dmela network data, we observed that the pink straight line (see Fig. 1 (right)) representing the power-law fit failed to align with many data points, particularly those corresponding to lower-degree nodes, indicating a poor fit in those regions. This suggests that when modeling with a single power-law, greater emphasis is placed on higher-degree nodes, while many lower-degree nodes, potentially containing valuable information about the network, are disregarded. Finally, our observation from Fig. 1 (right) reveals that the node degree distribution of the bio-Dmela network, when plotted on a log-log scale, exhibits nonlinearity and heavy-tailed nature upon closer inspection. This behavior differs from the pattern observed when a single power-law model is applied and has also been noted in many other biological networks (discussed in Sec V). Therefore, when fitting the degree distribution, it is crucial to consider all nodes, including both lower-degree and higher-degree nodes, to accurately capture the true pattern of the network’s degree distribution. This motivates the need to find a model that can accurately describe the main features of heavy-tailed and long-tailed behavior in the entire degree distribution of real-world complex networks. This article contributes to that effort by introducing a new heavy-tailed compounded Burr distribution. In the proposed model, the Poisson shift parameter ($\lambda$) gives compounded Burr distribution an advantage over static models like log-normal or standard power-law.
The proposed Poisson-shifted generalization and resulting CBurr model are not limited only to biological networks — they are applicable to a wide range of real-world complex networks (e.g., technological and social networks) that exhibit heavy-tailed degree distributions with finite-size effects or saturation behavior. The Poisson-shifting mechanism introduces a flexible exponential tail, allowing the model to capture non-monotonic behaviors and cutoff effects that arise naturally in complex real-world heavy-tailed networks. We focused on biological networks because they offer a biologically meaningful setting where both heavy-tailed degree distributions and saturation effects are prominently observed.